SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (May 2, 2006)

Attendees: Rainer Bartoldus, Al Eisner, Fang Fang, Selina Li, Rahmat Rahmat, Dong Su, Dan Walker

Operations

  • PEP's work on chamber replacement on Monday and Tuesday is going on smoothly. We should have beams delivered and start scrubbing this evening.
  • Chris O'Grady and Matt Weaver found and fixed another source of deadtime for EMT/GLT/IFR. See HN posted. Su Dong pointed out that the problem was probably already introduced when we started to have two ROMs (with slot 0 ROM added) running in these subsystems in 2001. We will be able to see its effect when we start taking physics data again.

EMT

  • Dan mentioned that we could try to change the EMT prescale factor. Rainer thinks that we might be okay with a prescale of 2. But for now, we will keep running with the factor being 3 and see the effect of the problem Chris and Matt fixed first.

DCT/DCZ

  • Selina created a web page to show the TSF firmware status. Rainer, Su Dong, Selina and Marc had a phone meeting last Thursday to discuss the different versions and the problems each version introduced. Marc will have more time to investigate after this Wednesday or Thursday.
  • The out-of-order and out-of-synch errors we have been seeing every once a while seem to correlate with the firmware used. The current version we use occasionally get the first 8 slots of the TSF-X boards out-of-order and out-of-synch from the last 8 slots. Su Dong thinks this is related to on-edge timing as the distance from the board to DCC matters. The TSF damages happen at a small enough rate that we can still tolerate till a final fix of the problem.
  • Fang saw ~10% logged events with ZPD damages as compared to the ~5% seen in L1. She will run L1TFMon on one of these runs with ZPD damage due to the TSF firmware. We suspected that the damages came during the time when the DCM lock was off and the board was trying to reset itself. See time history plots showing when the damages took place from the HN posted by Fang after the TRG operation meeting.

GLT

  • Su Dong will post something soon for the configuration study done with the 1Zn cut imposed.

Online

  • Fang updated the DCZ configuration status web page.
  • We are not going to have a new online release installed at IR2 yet since it's too disruptive to install a new offline release as it gets too close to the June 1st deadline for ICHEP.

L3

  • Rainer did a validation on the full derivatives and found that the L3 tracking efficiency increases by a few % but the reco efficiency decrease slightly. This study uses L1passthru runs processing withe new L3 tracking code. It turns out we are able to pick up a few more hadronic events, but loose some tracks too with the new code. The differences seen between the new and old L3 tracking codes still need to be understood first. The overall logging rate is not expected to get affected much. Rainer will post some of these results to the Trigger software HN.

Data Quality and Monitoring

  • There were three runs mislabeled on- or off-peak runs in run 5. See the first HN about this from last week. Al reminds us that the on- and off-peak boundary should be avoided for lumi ratio monitoring if we end up combining runs to get better statistics for monitoring. Rahmat would probably need a script to check if the run is on- or off-peak.
  • Rahmat didn't see anything wrong in OPR with the runs that contained ZPD damages mentioned last week. It looks like the ZPD damages have no impact on the ZPD efficiency. It's possible that the damages didn't affect data quality.
  • The driving force for the discussion of trickle injection monitoring comes from the proposal from PEP that the injection rate will be increased to 10+20 Hz as part of the plan for increasing luminosity. Al suggested to look at how lumi varies vs something in the Weaver plane. We probably need events inside the trickle inhibit window. Next week we will have a longer and deeper discussion of the trickle injection monitoring since we are running out of time in today's meeting.

This page is created by Selina Li (selina@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.